OpenStudy (anonymous):

Nancy left a bin outside in her garden to collect rainwater. She notices that 1 over 2 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words.

OpenStudy (anonymous):

@phi @mathmath333

OpenStudy (anonymous):

@imqwerty @dan915 @Elsa213 @JayJayTheWereWolfGal1

OpenStudy (phi):

does that say 1/2 gallons fills 2/3 of the barrel?

OpenStudy (anonymous):

Yes it does

OpenStudy (phi):

you could use ratios: 1/2 is to 2/3 as x is to 1 which says 1/2 gallon is to 2/3 of a bin as an unknown number of gallons (call it x) is to 1 (full) bin \[ \frac{ \frac{1}{2}}{\frac{2}{3}} = \frac{x}{1} \]

OpenStudy (phi):

or you could say 2/3 of a full bin is 1/2 gallon 2/3 x = 1/2 \[ \frac{2}{3} x= \frac{1}{2} \]

OpenStudy (anonymous):

OK the second one makes more sense

OpenStudy (phi):

to solve the second one, multiply both sides by 3/2 and simplify

OpenStudy (anonymous):

Multiply 12?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

you should get \[x= \frac{3}{2}\cdot \frac{1}{2} = \frac{3}{4} \] 3/4 of a gallon will fill the bin.